OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of f on the given interval. f(x) = (x^2 - 1)^3 [-1, 6] will fan and medal!!!

OpenStudy (freckles):

First step: Differentiate f Second step: Find critical numbers.

OpenStudy (freckles):

Third step: Is a matter of pluggin' into the original function

OpenStudy (aum):

Evaluate f(x) at the end points, x = -1 and x = 6 along with f(x) values at the critical point to determine the absolute max and min.

OpenStudy (anonymous):

f'(x)=(2x)^3 f'(x)=(6x)^2 @freckles

OpenStudy (freckles):

To differentiate f(x)=(x^2-1)^3 you need to chain rule, power rule, difference rule, power rule again, constant rule

OpenStudy (freckles):

\[f'(x)=3(x^2-1)^{3-1} \cdot (x^2-1)'\]

OpenStudy (anonymous):

to find the critical number, do I make f'(x) = 0 @freckles

OpenStudy (freckles):

you set f'=0 and solve for x have you finished finding f' yet?

OpenStudy (anonymous):

f'(x)= 6x^5-12x^3+6x

OpenStudy (freckles):

well I would have left it in factored form but okay

OpenStudy (freckles):

\[f'(x)=3(x^2-1)^{3-1}(x^2-1)' =3(x^2-1)^2(2x) \\ \text{ set} f'=0 \\ 3(x^2-1)^2(2x)=0 \text{ when } x=?\]

OpenStudy (anonymous):

x=0, x=1

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

there is x=-1 but that is endpoint anyways

OpenStudy (freckles):

now plug into f

OpenStudy (freckles):

plug the endpoints and the critical numbers which occured between the endpoints into f

OpenStudy (freckles):

\[f(-1)=? \\f(0)=? \\ f(1)=? \\f(6)=?\]

OpenStudy (anonymous):

f(-1)=-8 f(0)=-1 f(6)=42875 @freckles

OpenStudy (freckles):

how id you get -8 for the first one and then you find f(1) too?

OpenStudy (anonymous):

f(-1)=(-1^2)^3=-8 f(1)=0

OpenStudy (freckles):

\[f(-1)=((-1)^2-1)^3=(1-1)^3=0\]

OpenStudy (freckles):

both f(-1) and f(1) is zero

OpenStudy (freckles):

so the outputs we got were 0,-1,42875 - -1 is the lowest value (so that is the absolute min on the interval) 42875 is the highest value (so this is the absolute max on the interval)